OpenStudy (anonymous):

Given the literal equation ab/a+b = c , with b = –1 + a, what is the value of c if a = 5?

OpenStudy (anonymous):

A-20/9 B-25/9 C-4 D-11

OpenStudy (johnweldon1993):

okay so you have ab/a+b = c b = -1 + a so what a = 5...what is c? (5(-1 + 5))/(5 + (-1 +5)) distribute that 5 5(4)/ 5 + 4 20 / 9 hang on let me re-check that

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Find x if y = –4. 3y-6/x=12 A.0 B.-1/4 C.6 D.–4

OpenStudy (johnweldon1993):

wait hang on.....was 20/9 the answer? the only answer choice I see was -20/9???

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

20/9

OpenStudy (anonymous):

Find x if y = –4. 3y-6/x=12 A. 0 B. -1/4 C. 6 D. –4

OpenStudy (johnweldon1993):

\[\frac{ 3y - 6 }{ x } = 12\] or \[3y -\frac{ 6 }{ x } = 12\] which is it?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

or bottom one

OpenStudy (johnweldon1993):

okay so you have that equation you want to solve for x...and you know y = -4 3(-4) - (6/x) = 12 -12 - (6/x) = 12 add 12 to both sides here

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-24

OpenStudy (johnweldon1993):

-12 - (6/x) = 12 +12 +12 what does the equation look like now?

OpenStudy (anonymous):

6/x

OpenStudy (johnweldon1993):

-6/x = 24 now multiply both sides by -x 6 = -24x divide both sides by -24 x = -6/24 simplify this x = -1/4

OpenStudy (anonymous):

thank u

OpenStudy (johnweldon1993):

no problem